solve each literal equation for x.
pq + 2x/d = 4b

Answer by Theo(13342)

you want to solve for x.

this means you want to isolate x to the left side of the equation and have everything else to the right side of the equation.

start with pq + 2x/d = 4b

subtract pq from both sides of the equation to get:

2x/d = 4b - pq

multiply both sides of the equation by d to get:

2x = d * (4b - pq)

divide both sides of the equation by 2 to get:

x = (d * (4b - pq)) / 2

that's your solution.